# Quantum teleportation implies symmetry-protected topological order

### v2

@@ 3.1 Canonical form
Here we define the canonical form $\mathcal{W}$ of a measurement-based protocol $\mathcal{W}^{\prime}$. We note that physical unitaries $U$ that act after a dilated channel (i.e., $\mathcal{R}_{j}^{\prime}$ or $\mathcal{M}_{j}^{\prime}$) can always be “pulled through” to act before that channel, at the cost of modifying the latter according to $U\mathcal{R}_{j}^{\prime}=\mathcal{R}_{j}U$, where $\mathcal{R}_{j}=U\mathcal{R}_{j}^{\prime}U^{\dagger}$, and likewise for $\mathcal{M}_{j}^{\prime}$. See Fig. 3.
@@
We now briefly comment on Def. 4. Crucially, $\mathcal{W}$ and $\mathcal{W}^{\prime}$ are mathematically and physically the same protocol (i.e., $\mathcal{W}=\mathcal{W}^{\prime}$). However, the effective measured observables $A_{j}$ and recovery operators $\chi_{j}$ in $\mathcal{W}$ may differ from those in the naïve protocol $\mathcal{W}^{\prime}$. The canonical form (31) is essential to defining the class of protocols of interest and numerous other results. Additionally, familiar teleportation protocols (such as the cluster-state example in Sec. 2.4) already realize the canonical form (31).
However, by identifying modified channels as needed, any protocol $\mathcal{W}^{\prime}$ whose projective measurements are effectively instantaneous and whose recovery operations effectively realize a quantum circuit can be rewritten in canonical form $\mathcal{W}$ (31). The unitary operations captured by $\mathcal{U}$ may be generated by evolution under a local, time-dependent Hamiltonian. Using, e.g., the interaction picture, the corresponding unitaries can be “pulled through” the measurement channels and recovery operations without complication, because the latter are effectively instantaneous Friedman et al. (2022a). We note that, in the presence of outcome-dependent measurements, one also includes in $\mathcal{W}$ (31) a channel $\mathcal{M}^{\prime}$ between $\mathcal{M}$ and $\mathcal{R}$ comprising outcome-dependent measurements. However, such channels are generally not useful to physical teleportation (see Def. 3), and we do not consider them herein. Lastly, only protocols $\mathcal{W}^{\prime}=\mathcal{W}$ (31) that are already in canonical form saturate the teleportation bound (29) of Theorem 3.
Figure 3: (a) A naïve teleportation protocol $\mathcal{W}$ and (b) the equivalent canonical-form protocol $\mathcal{W}^{\prime}$ (31). The three layers of unitary gates are the same in both protocols; the measured observables $A_{j}^{\prime}$ (in red) are modified as the unitaries in their light cone (in orange) are “pulled through” to realize $A_{j}$ (32).
@@ 3.2 Measurement gates
The first restriction we impose relates to the projective measurements in the dilated channel $\mathcal{M}$ (16). In the literature on quantum information processing, by convention, one only considers single-qubit measurements corresponding to $Z_{j}$. From a theoretical standpoint, multiqubit projective measurements are universal Nielsen (2003); Leung (2002), and can achieve teleportation acting on a trivial (product) states; as a result, no meaningful constraints can be derived in their presence. From an experimental perspective, in most quantum hardware with qubits amenable to useful quantum tasks (such as teleportation), all measurements are implemented using a combination of physical unitaries and single-qubit $Z$ measurements (2.3).
Here, we restrict to the measurement of single-qubit observables $A_{j}^{\prime}$ in the naïve protocol $\mathcal{W}^{\prime}$. In the canonical-form protocol $\mathcal{W}$ (31), we allow observables $A_{j}$ that are unitarily equivalent to single-qubit operators, i.e.,
$A_{j}=U\,A_{j}^{\prime}\,U^{\dagger}\,,~{}~{}$ (32)
@@
The measurement of any nontrivial single-qubit observable $A_{j}$ with real eigenvalues $a_{0}>a_{1}$ is equivalent to measurement of its involutory part $\bar{A}_{j}$, defined to be
$\bar{A}_{j}$ $=\frac{1}{a_{0}-a_{1}}\sum\limits_{n=1}^{3}\,\operatorname*{tr}\limits_{j}\,\left[\,A_{j}\,\sigma^{n}_{j}\,\right]\,\sigma^{n}_{j}$ (33a)
@@
The proof of Prop. 4 appears in App. B. Importantly, Prop. 4 implies that all measurement channels take the simple form $\mathcal{M}_{j}$ (2.3), with $Z_{j}\to\bar{A}_{j}$ (4). If the effective observable is given by $A_{j}$ (32) in the canonical-form protocol $\mathcal{W}$ (31), then $\bar{A}_{j}=U\,\bar{A}_{j}^{\prime}U^{\dagger}$, where $\bar{A}_{j}^{\prime}$ is the involutory part of the single-qubit observable $A_{j}^{\prime}$ (unitary channels preserve involutions). As a result of Prop. 4, all measurements effectively have outcomes ±$\pm 1$, and spectral projectors $P^{(n)}_{A}=\left(\mathds{1}+(-1)^{n}\,\bar{A}\right)/2$ (15). This also simplifies the implementation of outcome-dependent operations, and facilitates the following result.
We next state Prop. 5, which establishes that, if the outcome of a measurement channel $\mathcal{M}_{j}$ (16) is not utilized later in the protocol $\mathcal{W}$ then, on average, $\mathcal{M}_{j}$ either acts trivially on all logical operators or causes $\mathcal{W}$ to fail.
@@
Consider the Heisenberg evolution of a logical operator $\Gamma\in\mathbf{P}_{*}{\rm ph}$ (2) through the channel $\mathcal{M}_{j}$ corresponding to the measurement of a single-qubit observable $A_{j}$ in the protocol $\mathcal{W}$, averaged over outcomes. If no aspect of $\mathcal{W}$ depends on the outcome of measuring $A_{j}$, then $\mathcal{M}_{j}$ either acts trivially on $\Gamma$ (i.e., $\Gamma\to\Gamma$), or trivializes $\Gamma$ (i.e., $\Gamma\to 0$).
@@ 3.3 Recovery operations
Like all useful quantum tasks, a teleportation protocol $\mathcal{W}$ outputs a state $\rho_{T}=\lvert\Psi_{T}\rangle\langle\Psi_{T}\rvert$ (7), from which, e.g., one extracts expectation values of observables. Hence, $\rho_{T}$ must be prepared and sampled many times; accordingly, when $\mathcal{W}$ involves measurements, the effective output state $\rho_{T}$ of $\mathcal{W}$ is the one that is prepared upon averaging over outcomes. When either $\rho_{T}$ or its logical part is pure, as with physical teleportation (see Def. 3), the above immediately implies the following Proposition.
@@
Suppose that a protocol $\mathcal{W}$ involving measurements and feedback outputs the state $\rho_{T}$ when averaged over outcomes. If the logical state $\rho_{\rm L}=\operatorname*{tr}_{F^{c}}\left[\rho_{T}\right]$ supported on $F\subset V$ is both pure and separable with respect to $F^{c}$, then $\mathcal{W}$ must realize $\rho_{\rm L}$ for every sequence of measurement outcomes.
The proof of Prop. 6 appears in App. D. Note that $F^{c}$ includes all Stinespring qubits, so that $\rho_{\rm L}$ is implicitly averaged over all outcomes. Additionally, Prop. 6 applies to protocols $\mathcal{W}$ for which $\rho_{T}$ itself is pure. As an aside, “virtual” teleportation protocols—in which the logical state is stored in the virtual legs of a tensor network Gross and Eisert (2007); Else et al. (2012)—generally do not realize the logical states $\lvert\psi_{n}\rangle$ on any finite subset $F\subset V$ of the physical qubits, even up to unitary operations on $F$. Instead, information is recovered from ostensibly unrelated operations Raussendorf et al. (2017); we relegate virtual teleportation to future work.
We now consider the implications for the outcome-dependent operations in $\mathcal{R}$. In general, using measurements to achieve a quantum task only outputs the desired state up to certain byproduct operators. The cost of using measurements to exceed the Lieb-Robinson bound (28) is that the state actually teleported is not $\lvert\psi\rangle$, but $\chi\,\lvert\psi\rangle$, where $\chi$ is the single-qubit byproduct operator, which is always invertible (i.e., unitary). For example, the cluster-state teleportation protocol considered in Sec. 2.4 outputs the logical state $Z^{n_{A}}_{f}\,X^{n_{B}}_{f}\,\lvert\psi\rangle$, rather than $\lvert\psi\rangle$ itself (on qubit $f$). Averaging over outcomes twirls the state $\rho_{\rm L}=\lvert\psi\rangle\langle\psi\rvert$ over the Pauli matrices, so that $\rho_{\rm L}=\mathds{1}/2$ is simply a random classical bit. The same conclusion is can be reached by considering the Heisenberg evolution of logical operators, as captured by Prop. 5.
At the same time, Prop. 6 implies that the errors introduced by measurements must be correctable for each sequence of outcomes $\boldsymbol{n}$. This is only possible if the measurement outcomes alone are sufficient to determine the byproduct operator $\chi$, and that operator is invertible, in which case one merely applies $\chi^{-1}$ to recover the logical state $\rho_{\rm L}$ for every $\boldsymbol{n}$. This is captured by the recovery channel $\mathcal{R}$ in $\mathcal{W}$ (31), which applies physical unitaries conditioned on the states of the Stinespring qubits, which are “read only” after the corresponding measurement.
@@
An outcome-dependent channel $\mathcal{R}_{\Omega}$ (18) is said to be Clifford if it can be written as
$\mathcal{R}_{\Omega}$ $=\mathds{1}\otimes\widetilde{P}^{(0)}_{\Omega}+\chi_{\Omega}\otimes\widetilde{P}^{(1)}_{\Omega}\,,~{}~{}$ (34)
in the canonical-form protocol $\mathcal{W}$ (31), up to multiplication on either side by elements of the physical Clifford group $\mathbf{C}_{*}{\rm ph}$, where $\chi_{\Omega}$ is a physical Pauli string (2), $\Omega\subset\widetilde{V}$ is a subset of the Stinespring qubits, and the operators
$\widetilde{P}^{(n)}_{\Omega}=\frac{1}{2}\Big{(}\widetilde{\mathds{1}}+(-1)^{n}\,\prod\limits_{j\in\Omega}\widetilde{Z}_{j}\Big{)}\,,~{}~{}$ (35)
project onto states of $\Omega$ with $\widetilde{Z}$ parity $(-1)^{n}$.
As far as we are aware, the recovery operations of all physical teleportation protocols known to the literature are Clifford in the sense of Def. 5. We also preclude outcome-dependent measurements, which are not only higher-order Clifford Gottesman and Chuang (1999) on $\mathcal{H}_{\rm dil}$ (14), but are generally not useful to physical teleportation on their own or in combination with the Clifford recovery channels of Def. 5. Essentially, conditional measurements fail to preserve crucial properties of the logical operators (e.g., unitarity).
As an aside, we conjecture that physical teleportation protocols can be classified according to their recovery group—the set generated by the recovery operators $\chi_{\Omega}$ for in the recovery channels $\mathcal{R}_{\Omega}$ (34) in $\mathcal{W}$ (31) under multiplication. That group is equivalent to the set of all possible byproduct operators. Restricting to the Clifford recovery channels of Def. 5, the recovery group must be a subset of the Pauli group $\mathbf{P}_{\rm ph}$. Since the phase of the logical states $\lvert\psi_{n}\rangle$ is unimportant, we have $XZ\cong ZX$, so that the recovery group is effectively Abelian. Since $\mathbf{P}_{\rm ph}$ also spans all physical operators, this is likely the simplest nontrivial recovery group that can realize. We relegate the consideration of more complicated recovery channels (and groups) to future work.
@@ 3.4 Standard teleportation protocols
A standard teleportation protocol $\mathcal{W}$ is a physical teleportation protocol (see Def. 3) for which the initial state $\lvert\Psi_{0}\rangle$ (6) is a product state and, when written in canonical form $\mathcal{W}=\mathcal{R}\,\mathcal{M}\,\mathcal{U}$ (31), (i) all measured observables $A_{j}$ are unitarily equivalent to single-qubit operators; (ii) all recovery operations are Clifford (34) in the sense of Def. 5; (iii) no unitaries are applied to sites already measured (in the Schrödinger picture); and (iv) the task distance $L=d(I,F)>vT$ (13) exceeds the Lieb-Robinson bound (28), where $v,T$ are determined by the local unitary $\mathcal{U}$, which is compatible with Theorem 2.
We now comment on the details standard teleportation (Def. 6), compared to the more general case of physical teleportation (Def. 3). We require that $L>vT$ to ensure that $\mathcal{W}$ is a teleportation—rather than state-transfer—protocol (see Def. 1). In combinatition with the requirement that $\lvert\Psi_{0}\rangle$ (6) be a product state, demanding $L>vT$ furnishes the locality-based analyses of Sec. 4, which also requires that $\lvert\Psi_{0}\rangle$ be separable with respect to a biparation of the graph $G$ compatible with Lemma 11. More generally, one might restrict to states $\lvert\Psi_{0}\rangle$ (6) that can be prepared using a FDQC $\mathcal{U}_{0}$; however, $\mathcal{U}_{0}$ can always be incorporated into $\mathcal{U}$ itself. The restriction to single-qubit measurements is motivated by experiment.
@@
The resource state $\lvert\Psi_{t}\rangle$ of a standard teleportation protocol $\mathcal{W}$ with canonical form $\mathcal{W}=\mathcal{R}\,\mathcal{M}\,\mathcal{U}$ (31) applied to the initial product state $\lvert\Psi_{0}(\boldsymbol{\alpha},\boldsymbol{\beta})\rangle$ (6) is
$\lvert\Psi_{t}\rangle=\mathcal{U}\,\lvert\Psi_{0}\rangle\,,~{}~{}$ (36)
where $\mathcal{U}$ realizes unitary time evolution on $\mathcal{H}_{\rm ph}$.
We note that the resource state $\lvert\Psi_{t}\rangle$ (36) is only defined on the sites on which $\mathcal{W}$ acts. For qubits on a graph $G$ in $d>1$ dimensions, a teleportation protocol $\mathcal{W}$ acts on some path $\cal C$ connecting the initial and final logical site(s), in which case the resource state $\lvert\Psi_{t}\rangle$ (36) refers only to the state of the qubits along the path $\mathcal{C}\subset G$.

### v3

@@ 3.1 Canonical form
Here we define the canonical form $\mathcal{W}$ of a measurement-based protocol $\mathcal{W}^{\prime}$. We note that physical unitaries $U$ that act after a dilated channel (i.e., $\mathcal{R}_{j}^{\prime}$ or $\mathcal{M}_{j}^{\prime}$) can always be “pulled through” to act before that channel, at the cost of modifying the latter according to $U\mathcal{R}_{j}^{\prime}=\mathcal{R}^{\vphantom{\prime}}_{j}U$, where $\mathcal{R}^{\vphantom{\prime}}_{j}=U\mathcal{R}_{j}^{\prime}U^{\dagger}$, and likewise for $\mathcal{M}_{j}^{\prime}$. See Fig. 3.
@@
We now briefly comment on Def. 4. Crucially, $\mathcal{W}$ and $\mathcal{W}^{\prime}$ are mathematically and physically the same protocol (i.e., $\mathcal{W}=\mathcal{W}^{\prime}$). However, the effective measured observables $A^{\,}_{j}$ and recovery operators $\chi^{\,}_{j}$ in $\mathcal{W}$ may differ from those in the naïve protocol $\mathcal{W}^{\prime}$. The canonical form (31) is essential to defining the class of protocols of interest and numerous other results. Additionally, familiar teleportation protocols (such as the cluster-state example in Sec. 2.4) already realize the canonical form (31).
However, by identifying modified channels as needed, any protocol $\mathcal{W}^{\prime}$ whose projective measurements are effectively instantaneous and whose recovery operations effectively realize a quantum circuit can be rewritten in canonical form $\mathcal{W}$ (31). The unitary operations captured by $\mathcal{U}$ may be generated by evolution under a local, time-dependent Hamiltonian. Using, e.g., the interaction picture, the corresponding unitaries can be “pulled through” the measurement channels and recovery operations without complication, because the latter are effectively instantaneous [42]. We note that, in the presence of outcome-dependent measurements, one also includes in $\mathcal{W}$ (31) a channel $\mathcal{M}^{\prime}$ between $\mathcal{M}$ and $\mathcal{R}$ comprising outcome-dependent measurements. However, such channels are generally not useful to physical teleportation (see Def. 3), and we do not consider them herein. Lastly, only protocols $\mathcal{W}^{\prime}=\mathcal{W}$ (31) that are already in canonical form saturate the teleportation bound (29) of Theorem 3.
Figure 3: (a) A naïve teleportation protocol $\mathcal{W}$ and (b) the equivalent canonical-form protocol $\mathcal{W}^{\prime}$ (31). The three layers of unitary gates are the same in both protocols; the measured observables $A_{j}^{\prime}$ (in red) are modified as the unitaries in their light cone (in orange) are “pulled through” to realize $A^{\,}_{j}$ (32).
@@ 3.2 Measurement gates
The first restriction we impose relates to the projective measurements in the dilated channel $\mathcal{M}$ (16). In the literature on quantum information processing, by convention, one only considers single-qubit measurements corresponding to $Z^{\vphantom{*}}_{j}$. From a theoretical standpoint, multiqubit projective measurements are universal [64, 65], and can achieve teleportation acting on a trivial (product) states; as a result, no meaningful constraints can be derived in their presence. From an experimental perspective, in most quantum hardware with qubits amenable to useful quantum tasks (such as teleportation), all measurements are implemented using a combination of physical unitaries and single-qubit $Z^{\vphantom{*}}$ measurements (2.3).
Here, we restrict to the measurement of single-qubit observables $A_{j}^{\prime}$ in the naïve protocol $\mathcal{W}^{\prime}$. In the canonical-form protocol $\mathcal{W}$ (31), we allow observables $A^{\,}_{j}$ that are unitarily equivalent to single-qubit operators, i.e.,
$A^{\vphantom{\prime}}_{j}=U\,A_{j}^{\prime}\,U^{\dagger}\,,~{}~{}$ (32)
@@
The measurement of any nontrivial single-qubit observable $A^{\,}_{j}$ with real eigenvalues $a^{\,}_{0}>a^{\,}_{1}$ is equivalent to measurement of its involutory part $\bar{A}^{\,}_{j}$, defined to be
$\bar{A}^{\vphantom{\dagger}}_{j}$ $=\frac{1}{a^{\,}_{0}-a^{\,}_{1}}\sum\limits_{n=1}^{3}\,\operatorname*{tr}\limits_{j}\,\left[\,A^{\vphantom{\prime}}_{j}\,\sigma^{n}_{j}\,\right]\,\sigma^{n}_{j}$ (33a)
@@
The proof of Prop. 4 appears in App. B. Importantly, Prop. 4 implies that all measurement channels take the simple form $\mathcal{M}^{\,}_{j}$ (2.3), with $Z^{\vphantom{*}}_{j}\to\bar{A}^{\,}_{j}$ (33). If the effective observable is given by $A^{\,}_{j}$ (32) in the canonical-form protocol $\mathcal{W}$ (31), then $\bar{A}^{\,}_{j}=U\,\bar{A}_{j}^{\prime}U^{\dagger}$, where $\bar{A}_{j}^{\prime}$ is the involutory part of the single-qubit observable $A_{j}^{\prime}$ (unitary channels preserve involutions). As a result of Prop. 4, all measurements effectively have outcomes ±$\pm 1$, and spectral projectors $P^{(n)}_{A}=\left(\mathds{1}+(-1)^{n}\,\bar{A}\right)/2$ (15). This also simplifies the implementation of outcome-dependent operations, and facilitates the following result.
We next state Prop. 5, which establishes that, if the outcome of a measurement channel $\mathcal{M}^{\,}_{j}$ (16) is not utilized later in the protocol $\mathcal{W}$ then, on average, $\mathcal{M}^{\,}_{j}$ either acts trivially on all logical operators or causes $\mathcal{W}$ to fail.
@@
Consider the Heisenberg evolution of a logical operator $\Gamma\in\mathbf{P}^{\vphantom{*}}_{*}{\rm ph}$ (2) through the channel $\mathcal{M}^{\,}_{j}$ corresponding to the measurement of a single-qubit observable $A^{\,}_{j}$ in the protocol $\mathcal{W}$, averaged over outcomes. If no aspect of $\mathcal{W}$ depends on the outcome of measuring $A^{\,}_{j}$, then $\mathcal{M}^{\,}_{j}$ either acts trivially on $\Gamma$ (i.e., $\Gamma\to\Gamma$), or trivializes $\Gamma$ (i.e., $\Gamma\to 0$).
@@ 3.3 Recovery operations
Like all useful quantum tasks, a teleportation protocol $\mathcal{W}$ outputs a state $\rho^{\,}_{T}=\lvert\Psi^{\,}_{T}\rangle\langle\Psi^{\,}_{T}\rvert$ (7), from which, e.g., one extracts expectation values of observables. Hence, $\rho^{\,}_{T}$ must be prepared and sampled many times; accordingly, when $\mathcal{W}$ involves measurements, the effective output state $\rho^{\,}_{T}$ of $\mathcal{W}$ is the one that is prepared upon averaging over outcomes. When either $\rho^{\,}_{T}$ or its logical part is pure, as with physical teleportation (see Def. 3), the above immediately implies the following Proposition.
@@
Suppose that a protocol $\mathcal{W}$ involving measurements and feedback outputs the state $\rho^{\,}_{T}$ when averaged over outcomes. If the logical state $\rho^{\,}_{\rm L}=\operatorname*{tr}_{F^{c}}\left[\rho^{\,}_{T}\right]$ supported on $F\subset V$ is both pure and separable with respect to $F^{c}$, then $\mathcal{W}$ must realize $\rho^{\,}_{\rm L}$ for every sequence of measurement outcomes.
The proof of Prop. 6 appears in App. D. Note that $F^{c}$ includes all Stinespring qubits, so that $\rho^{\,}_{\rm L}$ is implicitly averaged over all outcomes. Additionally, Prop. 6 applies to protocols $\mathcal{W}$ for which $\rho^{\,}_{T}$ itself is pure. As an aside, “virtual” teleportation protocols—in which the logical state is stored in the virtual legs of a tensor network [18, 23]—generally do not realize the logical states $\lvert\psi^{\,}_{n}\rangle$ on any finite subset $F\subset V$ of the physical qubits, even up to unitary operations on $F$. Instead, information is recovered from ostensibly unrelated operations [66]; we relegate virtual teleportation to future work.
We now consider the implications for the outcome-dependent operations in $\mathcal{R}$. In general, using measurements to achieve a quantum task only outputs the desired state up to certain byproduct operators. The cost of using measurements to exceed the Lieb-Robinson bound (28) is that the state actually teleported is not $\lvert\psi\rangle$, but $\chi\,\lvert\psi\rangle$, where $\chi$ is the single-qubit byproduct operator, which is always invertible (i.e., unitary). For example, the cluster-state teleportation protocol considered in Sec. 2.4 outputs the logical state $Z^{n^{\,}_{A}}_{f}\,X^{n^{\,}_{B}}_{f}\,\lvert\psi\rangle$, rather than $\lvert\psi\rangle$ itself (on qubit $f$). Averaging over outcomes twirls the state $\rho^{\,}_{\rm L}=\lvert\psi\rangle\langle\psi\rvert$ over the Pauli matrices, so that $\rho^{\,}_{\rm L}=\mathds{1}/2$ is simply a random classical bit. The same conclusion is can be reached by considering the Heisenberg evolution of logical operators, as captured by Prop. 5.
At the same time, Prop. 6 implies that the errors introduced by measurements must be correctable for each sequence of outcomes $\boldsymbol{n}$. This is only possible if the measurement outcomes alone are sufficient to determine the byproduct operator $\chi$, and that operator is invertible, in which case one merely applies $\chi^{-1}$ to recover the logical state $\rho^{\,}_{\rm L}$ for every $\boldsymbol{n}$. This is captured by the recovery channel $\mathcal{R}$ in $\mathcal{W}$ (31), which applies physical unitaries conditioned on the states of the Stinespring qubits, which are “read only” after the corresponding measurement.
@@
An outcome-dependent channel $\mathcal{R}^{\,}_{\Omega}$ (18) is said to be Clifford if it can be written as
$\mathcal{R}^{\vphantom{\dagger}}_{\Omega}$ $=\mathds{1}\otimes\widetilde{P}^{(0)}_{\Omega}+\chi^{\vphantom{\dagger}}_{\Omega}\otimes\widetilde{P}^{(1)}_{\Omega}\,,~{}~{}$ (34)
in the canonical-form protocol $\mathcal{W}$ (31), up to multiplication on either side by elements of the physical Clifford group $\mathbf{C}^{\vphantom{*}}_{*}{\rm ph}$, where $\chi^{\,}_{\Omega}$ is a physical Pauli string (2), $\Omega\subset\widetilde{V}$ is a subset of the Stinespring qubits, and the operators
$\widetilde{P}^{(n)}_{\Omega}=\frac{1}{2}\Big{(}\widetilde{\mathds{1}}^{\vphantom{*}}+(-1)^{n}\,\prod\limits_{j\in\Omega}\widetilde{Z}^{\vphantom{*}}_{j}\Big{)}\,,~{}~{}$ (35)
project onto states of $\Omega$ with $\widetilde{Z}^{\vphantom{*}}$ parity $(-1)^{n}$.
As far as we are aware, the recovery operations of all physical teleportation protocols known to the literature are Clifford in the sense of Def. 5. We also preclude outcome-dependent measurements, which are not only higher-order Clifford [3] on $\mathcal{H}^{\,}_{\rm dil}$ (14), but are generally not useful to physical teleportation on their own or in combination with the Clifford recovery channels of Def. 5. Essentially, conditional measurements fail to preserve crucial properties of the logical operators (e.g., unitarity).
As an aside, we conjecture that physical teleportation protocols can be classified according to their recovery group—the set generated by the recovery operators $\chi^{\,}_{\Omega}$ for in the recovery channels $\mathcal{R}^{\,}_{\Omega}$ (34) in $\mathcal{W}$ (31) under multiplication. That group is equivalent to the set of all possible byproduct operators. Restricting to the Clifford recovery channels of Def. 5, the recovery group must be a subset of the Pauli group $\mathbf{P}^{\vphantom{*}}_{\rm ph}$. Since the phase of the logical states $\lvert\psi^{\,}_{n}\rangle$ is unimportant, we have $X^{\vphantom{*}}Z^{\vphantom{*}}\cong Z^{\vphantom{*}}X^{\vphantom{*}}$, so that the recovery group is effectively Abelian. Since $\mathbf{P}^{\vphantom{*}}_{\rm ph}$ also spans all physical operators, this is likely the simplest nontrivial recovery group that can realize. We relegate the consideration of more complicated recovery channels (and groups) to future work.
@@ 3.4 Standard teleportation protocols
A standard teleportation protocol $\mathcal{W}$ is a physical teleportation protocol (see Def. 3) for which the initial state $\lvert\Psi^{\,}_{0}\rangle$ (6) is a product state and, when written in canonical form $\mathcal{W}=\mathcal{R}\,\mathcal{M}\,\mathcal{U}$ (31), (i) all measured observables $A^{\,}_{j}$ are unitarily equivalent to single-qubit operators; (ii) all recovery operations are Clifford (34) in the sense of Def. 5; (iii) no unitaries are applied to sites already measured (in the Schrödinger picture); and (iv) the task distance $L=d(I,F)>vT$ (13) exceeds the Lieb-Robinson bound (28), where $v,T$ are determined by the local unitary $\mathcal{U}$, which is compatible with Theorem 2.
We now comment on the details standard teleportation (Def. 6), compared to the more general case of physical teleportation (Def. 3). We require that $L>vT$ to ensure that $\mathcal{W}$ is a teleportation—rather than state-transfer—protocol (see Def. 1). In combinatition with the requirement that $\lvert\Psi^{\,}_{0}\rangle$ (6) be a product state, demanding $L>vT$ furnishes the locality-based analyses of Sec. 4, which also requires that $\lvert\Psi^{\,}_{0}\rangle$ be separable with respect to a biparation of the graph $G$ compatible with Lemma 11. More generally, one might restrict to states $\lvert\Psi^{\,}_{0}\rangle$ (6) that can be prepared using a FDQC $\mathcal{U}^{\,}_{0}$; however, $\mathcal{U}^{\,}_{0}$ can always be incorporated into $\mathcal{U}$ itself. The restriction to single-qubit measurements is motivated by experiment.
@@
The resource state $\lvert\Psi^{\,}_{t}\rangle$ of a standard teleportation protocol $\mathcal{W}$ with canonical form $\mathcal{W}=\mathcal{R}\,\mathcal{M}\,\mathcal{U}$ (31) applied to the initial product state $\lvert\Psi^{\,}_{0}(\boldsymbol{\alpha},\boldsymbol{\beta})\rangle$ (6) is
$\lvert\Psi^{\vphantom{\prime}}_{t}\rangle=\mathcal{U}\,\lvert\Psi^{\vphantom{\prime}}_{0}\rangle\,,~{}~{}$ (36)
where $\mathcal{U}$ realizes unitary time evolution on $\mathcal{H}^{\,}_{\rm ph}$.
We note that the resource state $\lvert\Psi^{\,}_{t}\rangle$ (36) is only defined on the sites on which $\mathcal{W}$ acts. For qubits on a graph $G$ in $d>1$ dimensions, a teleportation protocol $\mathcal{W}$ acts on some path $\cal C$ connecting the initial and final logical site(s), in which case the resource state $\lvert\Psi^{\,}_{t}\rangle$ (36) refers only to the state of the qubits along the path $\mathcal{C}\subset G$.